OpenStudy (anonymous):

solve 3x^2 - 2x -9 =0 by completing the square

OpenStudy (anonymous):

please help me :)

OpenStudy (whpalmer4):

to complete the square, you want to arrange your equation to look like \[x^2 + bx = c\] I'd start out by adding 9 to each side, and then dividing each side by 3

OpenStudy (anonymous):

ok, i did that so i would have x^2-2/3x+3=9 ?

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

[3x^2-2x-9=0]/3 (divide everything by 3) x^2 - (2/3)x -3 = 0 x^2 -(2/3)x = 3 x^2 - (2/3)x + (1/3)^2 = 3 + (1/3)^2 (x-1/3)^2 = 28/9 sqrt((x-1/3)^2) = +- sqrt(28/9) x-1/3 = +- (2sqrt(7))/3 x = (1/3)+- (2sqrt(7)/3) or x = [1+- 2sqrt(7)]/3 I hope you get what I wrote.

OpenStudy (anonymous):

i mean x^2-2/3x=3...

OpenStudy (anonymous):

why did you had (1/3)^2 to each side???

OpenStudy (anonymous):

*add

OpenStudy (whpalmer4):

\[x^2 - \frac{2}{3}x =3\] Now we take half of the coefficient of \(x\), giving us \(-\frac{1}{3}\) we square that and add to both sides \[x^2 -\frac{2}{3}x + (-\frac{1}{3})^2 = 3 + (-\frac{1}{3})^2\]

OpenStudy (anonymous):

had to add that to complete the square on left side. and whatever you add on left you should do it to right too. (2/3)/2 = 1/3 then you square it to complete the square

OpenStudy (whpalmer4):

now we can rewrite the left side as \[(x-\frac{1}{3})^2 = 3 + (-\frac{1}{3})^2\]

OpenStudy (whpalmer4):

we have to add the same thing to each side or we break the equation. \[a = a\]\[a+b = a+b\]

OpenStudy (anonymous):

oooohhh I see

OpenStudy (anonymous):

so whats the next step?

OpenStudy (whpalmer4):

if we multiply \[(x+a)(x+a) = x^2+ ax + ax + a^2 = x^2 + 2ax + a^2\]you see where the "take half of the coefficient of the x term and square it" comes from.

OpenStudy (anonymous):

ok...

OpenStudy (whpalmer4):

if we pretend the coefficient of the x term \(= 2a\) then we can rewrite the left hand side as \((x + a)^2\). that's where the term "completing the square" comes from.

OpenStudy (anonymous):

is it x-1/3= 3+/- √-1/3 ???

OpenStudy (anonymous):

next will be: the left side is now a complete square \[x ^{2}-\frac{ 2 }{ 3 }x + (\frac{ 1 }{ 3 })^{2} =\frac{ 28 }{ 9 }\]

OpenStudy (anonymous):

wait, how did you get that??

OpenStudy (whpalmer4):

\[(x-\frac{1}{3})^2 = 3 + (-\frac{1}{3})^2\] \[(x-\frac{1}{3})^2 = 3*\frac{9}9 + (-\frac{1}{3})^2 = \frac{27}{9}+\frac{1}{9} = \frac{28}{9}\] take square root of both sides: \[(x-\frac{1}{3}) =\pm \sqrt{\frac{28}{9}} =\pm \frac{2}{3}\sqrt{7}\]

OpenStudy (anonymous):

i like this format, thank you!

OpenStudy (anonymous):

ok I have a couple ?s... where did you get the 9/9 from?

OpenStudy (whpalmer4):

9/9 = 1, right?

OpenStudy (whpalmer4):

I'm just making a common denominator, because (1/3)^2 = 1/9

OpenStudy (anonymous):

yes, but why 9/9

OpenStudy (anonymous):

oh

OpenStudy (whpalmer4):

converting 3 to n/9 27/9 = 3, right?

OpenStudy (anonymous):

yes, so my final answer is x-1/3= +/- 2/3√7 ?

OpenStudy (anonymous):

I'm sorry, this just doesn't make sense? the x has to be by itself right?

OpenStudy (anonymous):

move that -1/3 to left side and it will become positive 1/3

OpenStudy (anonymous):

you mean the right side?

OpenStudy (anonymous):

\[x= \frac{ 1 }{ 3 } \pm \frac{ 2\sqrt{7} }{ 3}\]

OpenStudy (anonymous):

in short you are adding 1/3 to both side of equation.

OpenStudy (anonymous):

ok, now can u please explain to me how u got the 2√7/3 part?

OpenStudy (whpalmer4):

\[\sqrt{\frac{28}{9}} = \sqrt{\frac{2*2*7}{3*3}} =\frac{2}{3}\sqrt{7}\]

OpenStudy (anonymous):

ok: \[x= \frac{ 1 }{ 3 }\pm \frac{ \sqrt{28} }{ 3 }\] then \[\sqrt{28} = \sqrt{4\times7}\] right? so we take out 4 since \[\sqrt{4} = 2\] results to: \[\sqrt{4\times7} = 2\sqrt{7}\] therefore, the equation is: \[x = \frac{ 1 }{ 3 }\pm \frac{ 2\sqrt{7} }{ 3 }\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

I understand... you guys are great! Thank you both :)

OpenStudy (anonymous):

would u be able to help me with one more?